# Size-induced Exchange Bias in Single-phase CoO Nanoparticles

### v1

@@ 1 Introduction
Exchange bias refers to the shift of the magnetic hysteresis loop of a material along the field axis 1, 2, 3, 4. It appears in a system with two different exchange-coupled magnetic phases after cooling in the presence of an external magnetic field. This phenomenon is well-known as conventional exchange bias (CEB) and has been observed in heterostructures consisting of ferromagnetic (FM)-antiferromagnetic (AFM), FM-spin glass, AFM-ferrimagnetic, FM-ferrimagnetic, and FM-FM phases 1, 2, 3, 4, 5, 6, 7, 8. Besides heterostructures, CEB has also been reported in nanoparticles due to naturally occurring two magnetic phases, predominantly with an antiferromagnetic ground state where surface spins play a significant role 9, 10, 11, 12, 13, 14, 15. Notably, a prominent exchange bias effect even in the ZFC condition, known as spontaneous exchange bias (SEB), has been reported in a few systems 16, 17, 18, 19. Until now, SEB has been reported in systems such as IrMn/FeCo bilayers, Ni-Mn-Sn alloys, Si-substituted MnNiSnSi system, BiFeO-BiFeO nanocomposites 2, 3, 6, 20, 21, 22. The exchange bias effect plays a pivotal role in both fundamental research and numerous applications, including permanent magnets, magnetic resonance imaging, and spintronic devices such as magnetic tunnel junctions, magnetoresistance sensors, and magnetic random-access memory 23, 24, 25.
In this context, AFM nanoparticles could be intriguing, given the potential occurrence of two magnetic phases naturally with decrease in their size 10, 11, 26. The magnetic properties of CoO nanoparticles are less explored compared to other oxide nanostructures like NiO since their synthesis is challenging due to the greater thermodynamic stability of the CoO phase, leading to the immediate reducibility of CoO to Co metal 27. Furthermore, there is no report on SEB and CEB effects in single-phase nanosized CoO, where two exchange-coupled magnetic phases occur naturally with the decrease in crystallite size 28, 29, 30. This lack of information motivated us to explore the size-dependent magnetic properties of CoO nanoparticles further. Recently, we developed a method for synthesizing separate single-phase face-centered cubic (fcc) and monoclinic phases of reasonably monodispersed CoO nanoparticles with varying crystallite size 27. The dominant role of trioctylphosphine (TOP) was found to influence the variation in crystallite size during synthesis. The development of new methods for the synthesis of magnetic nanostructures is of particular interest as they have many technological applications 31, 32. For example, Chen et al. demonstrated a significant enhancement in the charge transfer process of BiVO/CoO via Fe doping, which is useful for water splitting 33.
@@ 2 Experimental details
In the present study, we utilized well-characterized single-phase CoO nanoparticles, Co/CoO nanocomposites, and silica-coated CoO nanoparticles, denoted as CoO@SiO27. These samples were prepared using thermal decomposition method, and the crystallite size was varied by changing the amount of trioctylphosphine (TOP)27. Typically, 0.25 ml of preheated TOP at 200 C was added to a solution of 2 g of Co(CHCOO)·4HO and 8 ml of oleylamine (OA), already degassed at 120–130 C for half an hour. The resulting solution was further heated to 230 C for 1.5 h under an argon atmosphere. After cooling down to 30 C, the reaction product was washed and dried; the sample was denoted as CoO4. Other samples, coded as CoO6 and CoO8, were prepared with TOP concentrations of 3 and 12 ml, respectively, with other conditions remaining the same. Additionally, we prepared an extra sample, CoO10, of Co/CoO nanocomposites with a combination of 8 ml of OA and 12 ml of TOP at 260 C in 2 hrs, with other conditions remaining constant. This nanocomposite sample of Co/CoO was prepared to compare the magnetic properties with single-phase CoO nanoparticles since Co/CoO is a prototype example of the exchange bias effect, which is the focus of this paper.
Furthermore, we coated SiO on the prepared powder of the smallest-sized nanoparticle sample CoO8 with the specific aim of observing the influence of surface spins on magnetic properties. To functionalize SiO on the surface of CoO8 nanoparticles, we employed a standard Stöber method described elsewhere 34. In brief, 0.10 g of 10.8 nm nanoparticles were ultrasonicated in a mixture of 40 ml of ethanol, 10 ml of deionized water (DIW), and 1.2 ml of 28 wt. ammonia solution for 1 h. Then, 0.43 mL of tetraethyl orthosilicate (TEOS) was added dropwise to this mixture. After stirring for 6 h, the obtained product was separated by centrifugation for 15 min at 9000 rpm and then washed with DIW. The product was dried at 100 °C and then crushed well to make a fine powder.
@@ 3.1 Structural study
The crystal structure of CoO3, CoO6, and CoO8 was identified using XRD patterns. Figures 1 (a-c) display the laboratory XRD patterns along with the Rietveld refinement of CoO3, CoO6, and CoO8. The five peaks appearing around 36.3, 42.6, 61.4, 73.5, and 77.4 in the XRD pattern of CoO3 correspond to (1 1 1), (2 0 0), (2 2 0), (3 1 1), and (2 2 2) planes, respectively, of the fcc phase of CoO (Fig. 1 (a)). The similar five peaks in the XRD patterns of CoO6 and CoO8 confirm their fcc phase (Figs. 1 (b, c)). Notably, the peaks are broadened from CoO3 to CoO8, indicating a reduction in crystallite size with an increase in the concentration of TOP 27, 35, 36. The average crystallite size was evaluated using the Debye-Scherrer formula, defined as = , where is the crystallite size, is the full width at half maximum (FWHM) of the peak, represents the Bragg angle, and is an anisotropic constant in the range of 0.87–1.0. Here, we used = 0.9 considering the spherical nanoparticles. The crystallite size is found to be 34.60.8 nm, 22.90.7 nm, and 10.80.9 nm for CoO3, CoO6, and CoO8, respectively, listed in Table 1. Furthermore, the XRD patterns were analyzed using the Rietveld refinement method implemented with the FullProf Suite package 37. Rietveld analysis reveals that the prepared nanoparticle samples of CoO crystallize in the rock-salt cubic or face-centered cubic (fcc) phase with space group Fm-3m (No. 225). The obtained lattice parameters are a = 4.267, 4.265, and 4.258 Å for CoO3, CoO6, and CoO8, respectively. Although these samples, CoO3, CoO6, and CoO8, consist of single-phase CoO nanoparticles, their magnetic properties vary significantly due to distinct crystallite sizes, primarily influenced by the dominant role of finite size and surface effects. As crystallite size decreases, surface effects become more prominent, leading to different magnetic behavior at the surface compared to the core of the nanoparticles, as discussed later.
@@
We have also performed synchrotron radiation XRD of CoO3, CoO6, CoO8, and CoO10 (Fig. 2 (a)). The four peaks around 36.4, 42.3, 61.4, and 73.6 correspond to (1 1 1), (2 0 0), (2 2 0), and (3 1 1) planes, respectively, for CoO3, CoO6, and CoO8 have been observed. However, the XRD pattern of CoO10 shows an extra peak relative to CoO10, corresponding to Co, indicating that this sample is a composite of Co and CoO. The progressive increase in peak width observed in XRD peaks from CoO3 to CoO8 signifies a corresponding decrease in crystallite size, consistent with findings from the Lab XRD.
@@
The crystal structure of rock-salt cubic CoO is visualized using VESTA 38 as shown in (Fig. 2 (b)). The crystal structure shows that the effective number of Co atoms is four per unit cell, located at the corners and face centers of the unit cell. The Co is bonded to six equivalent O atoms to form a mixture of edge and corner-sharing CoO6 octahedra. The Co–O bond lengths are about 2.14 . The Co and O atoms have only one site with 4 and 4 Wyckoff positions, respectively.
@@ 3.2 X-ray absorption near edge spectroscopy
We employed XANES to determine the valence state of the Co ion in the prepared samples. We present the normalized XANE spectra of Co 2p and O 1s for CoO3, CoO6, and CoO8, along with bulk CoO as shown in Fig. 3 (a). The Co -edge splits into two main features due to core level spin-orbit coupling, denoted as the lower energy and the higher energy , which are assigned to the transitions from Co 2p () and 2p () electrons into the empty 3d states, respectively.The energy separation between and edges in CoO3 is found to be about 15.7 eV, well-matched with the value obtained from photoelectron spectroscopy (XPS)27. Furthermore, the Co edge is split into four features at around 776.1 (A), 777.6 (B), 778.8 (C), 781.0 (D), mainly due to electron-electron interactions39. Therefore, the Co L edge of CoO3 exhibits five main features near 776.1 (A), 777.6 (B), 778.8 (C), 781.0 (D), and 793.4 (E) eV, between 772.5 and 798.5 eV, which are in agreement with earlier reports 39, 40. CoO6 and CoO8 also show similar five features. The XANE spectrum of CoO shows mainly three peaks around 777.7, 779.4, and 794.0 eV, in line with an earlier report 39. The Co edge of CoO has only two main features. Note that the multiplet structure of the -edge is highly sensitive to the coordination of the Co ion. The Co 2p spectrum of CoO is due to octahedral Co crystal field splitting and is significantly different from CoO, in which both octahedral Co and tetrahedral Co coordinations are present.
@@
The 2p orbitals of the oxygen ligand are involved in the bonding configuration with Co metal ions. Therefore, dipole-allowed 1s to 2p transitions of O 1s can directly provide information about the oxygen charge state and reflect the partial density of O 2p unoccupied electronic states in the conduction band 41, 42. We measured the O 1s XANE spectra of CoO3, CoO6, and CoO8 along with bulk CoO, as shown in Fig. 3(b).The O 1s spectra of these samples exhibit mainly five features at around 529.5 (P1), 532.3 (P2), 536.5 (P3), 539.8 (P4), and 549.5 (P5) eV, ranging between 525.5 to 552.5 eV. The peaks P2 to P5 in these nanoparticles are in close agreement with earlier reports 42, 39, 40. However, an additional pre-edge feature, P1, is found compared to bulk CoO, and it becomes more pronounced with the decrease in crystallite size. This feature indicates that the electronic structure is modified in nanoparticles compared to the bulk counterpart. The features between 530 and 535 eV correspond to transitions to unoccupied electronic states of O 2p character hybridized with Co 3d states 42. Above 535 eV, features in XANE are mainly associated with transitions to O 2p character hybridized with Co 4s, 4p states 42. The O 1s character of CoO is significantly different from that of CoO nanoparticles. For example, a sharp peak that appears in CoO is absent in CoO nanoparticles. These results indicate that the bonding configuration of O ions with Co metal ions, and also valence states, are different in CoO nanoparticles and CoO. Therefore, the XANES results of CoO3, CoO6, and CoO8 clearly show that these nanoparticles exhibit a pure Co valence state, which is in line with XRD data
@@ 3.3 Magnetic Properties
We have also recorded the loop at 5 K in the FC condition to probe the conventional exchange bias (Fig.6(b, c)). The sample has been cooled to 5 K from 400 K in the presence of = 7 T, and the field is cycled between 7 T, which is sufficient to obtain technical saturation and hence rules out the possibility of the minor hysteresis loop. In this case, we have recorded the loops of only CoO6 and CoO8, which show prominent hysteresis at 5 K in ZFC (Fig.5(b, c)). Under the FC condition, an enhancement in coercivity as compared to the ZFC protocol has been found in the CoO nanoparticles (Fig.6(b, c)). The hysteresis loop of CoO8 at 5 K exhibits a larger EB of around Oe in FC at 5 K. In contrast, although CoO6 does not show any exchange bias in the ZFC condition, it shows Oe in FC at 5 K. In this context, we note the reports of Lierop et al. of SEB and CEB of around 75 and 150 Oe, respectively, in NiFe/CoO thin film21, and Takano et al. estimated and observed CEB of Oe on NiFe/CoO bilayers 8. The value of CEB in the CoO nanoparticles is larger compared to earlier reports on NiFe/CoO bilayers and NiFe/CoO thin film 21, 8.
@@
(ii) CoO8, i.e., 10.8 nm CoO nanoparticles coated with diamagnetic silica (SiO).
@@
When we coated CoO8 with SiO, which is supposed to have the same crystallite size, it influences the magnetic response of nanoparticles quite interestingly. It should be mentioned here that although the thickness of the silica coating and hence the mass of the CoO/SiO system is not precisely known, its thickness is believed to be smaller than the diameter of the nanoparticle 34. The temperature-dependent magnetization measured in ZFC and FC at 100 Oe is shown in Fig.9(a). As temperature decreases, ZFC magnetization decreases slowly compared to CoO8 (Fig.4(c)) and shows a steep increase below 20 K (Fig. 9(a), inset). However, a clear hump associated with is seen near the same temperature of 213 K as found also in CoO8. This means SiO coating just affects the surface moment but not the core of the nanoparticles. The magnetization measured in FC at 7 T is approximately independent of temperature down to 20 K, and then steeply increases as temperature further decreases, similar to ZFC magnetization.
The curve at 300 K of SiO-coated CoO nanoparticles is depicted in Fig.9(b). It shows finite coercivity and remanence, although it does not saturate till Tesla, which indicates a weak ferromagnetic response of CoO8 still present after SiO coating. The value of the coercive field has been found to be around Oe, which is significantly reduced compared to CoO8 (610 Oe). Note that both the and curves of SiO-coated CoO nanoparticles show distinct behavior compared to bare CoO8, which underlines the fact that SiO coating affects the overall magnetic response of the CoO nanoparticles. SiO-coating on nanoparticles is expected to reduce agglomeration of nanoparticles and likely to reduce the interparticle interaction43, 51. Thus, the SiO-coating may affect the overall collective magnetic response. In addition, the magnetism at the interface between magnetic CoO nanoparticles and non-magnetic SiO can be quite interesting and is a subject of detailed investigation 52, 53.
Therefore, it is noteworthy to mention a few distinct changes in the magnetic response of Co/CoO nanocomposites and SiO-coated CoO8 nanoparticles compared to bare CoO8 nanoparticles:
@@
(ii) The magnetization of CoO8 is significantly reduced after SiO coating. Since the magnetic moment in CoO8 mainly comes from surface spins, the reduction in the magnetic moment may indicate that surface spins are affected by SiO.
(iii) The is observed near the same temperature of 213 K. This suggests that SiO coating mainly affects the surface moment but not the core of nanoparticles.
(iv) The bifurcation between ZFC and FC curves has reduced. This indicates that spin disorder reduces after SiO coating, possibly originating from the shell of nanoparticles.
(v) A steep rise in magnetization below 20 K, similar to the larger-sized sample CoO3, is observed. This is consistent with the reduction in magnetization with SiO, similar to an increase in size as in CoO3.

### v2

@@ 1 Introduction
Exchange bias refers to the shift of the magnetic hysteresis loop of a material along the field axis 1, 2, 3, 4. It appears in a system with two different exchange-coupled magnetic phases after cooling in the presence of an external magnetic field. This phenomenon is well-known as conventional exchange bias (CEB) and has been observed in heterostructures consisting of ferromagnetic (FM)-antiferromagnetic (AFM), FM-spin glass, AFM-ferrimagnetic, FM-ferrimagnetic, and FM-FM phases 1, 2, 3, 4, 5, 6, 7, 8. Besides heterostructures, CEB has also been reported in nanoparticles due to naturally occurring two magnetic phases, predominantly with an antiferromagnetic ground state where surface spins play a significant role 9, 10, 11, 12, 13, 14, 15. Notably, a prominent exchange bias effect even in the ZFC condition, known as spontaneous exchange bias (SEB), has been reported in a few systems 16, 17, 18, 19. Until now, SEB has been reported in systems such as IrMn/FeCo bilayers, Ni-Mn-Sn alloys, Si-substituted MnNiSnSi system, BiFeO3-Bi2Fe4O9 nanocomposites 2, 3, 6, 20, 21, 22. The exchange bias effect plays a pivotal role in both fundamental research and numerous applications, including permanent magnets, magnetic resonance imaging, and spintronic devices such as magnetic tunnel junctions, magnetoresistance sensors, and magnetic random-access memory 23, 24, 25.
In this context, AFM nanoparticles could be intriguing, given the potential occurrence of two magnetic phases naturally with decrease in their size 10, 11, 26. The magnetic properties of CoO nanoparticles are less explored compared to other oxide nanostructures like NiO since their synthesis is challenging due to the greater thermodynamic stability of the Co3O4 phase, leading to the immediate reducibility of CoO to Co metal 27. Furthermore, there is no report on SEB and CEB effects in single-phase nanosized CoO, where two exchange-coupled magnetic phases occur naturally with the decrease in crystallite size 28, 29, 30. This lack of information motivated us to explore the size-dependent magnetic properties of CoO nanoparticles further. Recently, we developed a method for synthesizing separate single-phase face-centered cubic (fcc) and monoclinic phases of reasonably monodispersed CoO nanoparticles with varying crystallite size 27. The dominant role of trioctylphosphine (TOP) was found to influence the variation in crystallite size during synthesis. The development of new methods for the synthesis of magnetic nanostructures is of particular interest as they have many technological applications 31, 32. For example, Chen et al. demonstrated a significant enhancement in the charge transfer process of BiVO4/CoOx via Fe doping, which is useful for water splitting 33.
@@ 2 Experimental details
In the present study, we utilized well-characterized single-phase CoO nanoparticles, Co/CoO nanocomposites, and silica-coated CoO nanoparticles, denoted as CoO@SiO227. These samples were prepared using thermal decomposition method, and the crystallite size was varied by changing the amount of trioctylphosphine (TOP)27. Typically, 0.25 ml of preheated TOP at 200 ∘C was added to a solution of 2 g of Co(CH3COO)2·4H2O and 8 ml of oleylamine (OA), already degassed at 120–130 ∘C for half an hour. The resulting solution was further heated to 230 ∘C for 1.5 h under an argon atmosphere. After cooling down to 30 ∘C, the reaction product was washed and dried; the sample was denoted as CoO4. Other samples, coded as CoO6 and CoO8, were prepared with TOP concentrations of 3 and 12 ml, respectively, with other conditions remaining the same. Additionally, we prepared an extra sample, CoO10, of Co/CoO nanocomposites with a combination of 8 ml of OA and 12 ml of TOP at 260 ∘C in 2 hrs, with other conditions remaining constant. This nanocomposite sample of Co/CoO was prepared to compare the magnetic properties with single-phase CoO nanoparticles since Co/CoO is a prototype example of the exchange bias effect, which is the focus of this paper.
Furthermore, we coated SiO2 on the prepared powder of the smallest-sized nanoparticle sample CoO8 with the specific aim of observing the influence of surface spins on magnetic properties. To functionalize SiO2 on the surface of CoO8 nanoparticles, we employed a standard Stöber method described elsewhere 34. In brief, 0.10 g of 10.8 nm nanoparticles were ultrasonicated in a mixture of 40 ml of ethanol, 10 ml of deionized water (DIW), and 1.2 ml of 28 wt. ammonia solution for 1 h. Then, 0.43 mL of tetraethyl orthosilicate (TEOS) was added dropwise to this mixture. After stirring for 6 h, the obtained product was separated by centrifugation for 15 min at 9000 rpm and then washed with DIW. The product was dried at 100 °C and then crushed well to make a fine powder.
@@ 3.1 Structural study
The crystal structure of CoO3, CoO6, and CoO8 was identified using XRD patterns. Figures 1 (a-c) display the laboratory XRD patterns along with the Rietveld refinement of CoO3, CoO6, and CoO8. The five peaks appearing around 36.3∘, 42.6∘, 61.4∘, 73.5∘, and 77.4∘ in the XRD pattern of CoO3 correspond to (1 1 1), (2 0 0), (2 2 0), (3 1 1), and (2 2 2) planes, respectively, of the fcc phase of CoO (Fig. 1 (a)). The similar five peaks in the XRD patterns of CoO6 and CoO8 confirm their fcc phase (Figs. 1 (b, c)). Notably, the peaks are broadened from CoO3 to CoO8, indicating a reduction in crystallite size with an increase in the concentration of TOP 27, 35, 36. The average crystallite size was evaluated using the Debye-Scherrer formula, defined as = , where is the crystallite size, is the full width at half maximum (FWHM) of the peak, represents the Bragg angle, and is an anisotropic constant in the range of 0.87–1.0. Here, we used = 0.9 considering the spherical nanoparticles. The crystallite size is found to be 34.60.8 nm, 22.90.7 nm, and 10.80.9 nm for CoO3, CoO6, and CoO8, respectively, listed in Table 1. Furthermore, the XRD patterns were analyzed using the Rietveld refinement method implemented with the FullProf Suite package 37. Rietveld analysis reveals that the prepared nanoparticle samples of CoO crystallize in the rock-salt cubic or face-centered cubic (fcc) phase with space group Fm-3m (No. 225). The obtained lattice parameters are a = 4.267, 4.265, and 4.258 Å for CoO3, CoO6, and CoO8, respectively. Although these samples, CoO3, CoO6, and CoO8, consist of single-phase CoO nanoparticles, their magnetic properties vary significantly due to distinct crystallite sizes, primarily influenced by the dominant role of finite size and surface effects. As crystallite size decreases, surface effects become more prominent, leading to different magnetic behavior at the surface compared to the core of the nanoparticles, as discussed later.
@@
We have also performed synchrotron radiation XRD of CoO3, CoO6, CoO8, and CoO10 (Fig. 2 (a)). The four peaks around 36.4∘, 42.3∘, 61.4∘, and 73.6∘ correspond to (1 1 1), (2 0 0), (2 2 0), and (3 1 1) planes, respectively, for CoO3, CoO6, and CoO8 have been observed. However, the XRD pattern of CoO10 shows an extra peak relative to CoO10, corresponding to Co, indicating that this sample is a composite of Co and CoO. The progressive increase in peak width observed in XRD peaks from CoO3 to CoO8 signifies a corresponding decrease in crystallite size, consistent with findings from the Lab XRD.
@@
The crystal structure of rock-salt cubic CoO is visualized using VESTA 38 as shown in (Fig. 2 (b)). The crystal structure shows that the effective number of Co atoms is four per unit cell, located at the corners and face centers of the unit cell. The Co2+ is bonded to six equivalent O2- atoms to form a mixture of edge and corner-sharing CoO6 octahedra. The Co–O bond lengths are about 2.14 . The Co and O atoms have only one site with 4 and 4 Wyckoff positions, respectively.
@@ 3.2 X-ray absorption near edge spectroscopy
We employed XANES to determine the valence state of the Co ion in the prepared samples. We present the normalized XANE spectra of Co 2p and O 1s for CoO3, CoO6, and CoO8, along with bulk Co3O4 as shown in Fig. 3 (a). The Co -edge splits into two main features due to core level spin-orbit coupling, denoted as the lower energy and the higher energy , which are assigned to the transitions from Co 2p () and 2p () electrons into the empty 3d states, respectively.The energy separation between and edges in CoO3 is found to be about 15.7 eV, well-matched with the value obtained from photoelectron spectroscopy (XPS)27. Furthermore, the Co edge is split into four features at around 776.1 (A), 777.6 (B), 778.8 (C), 781.0 (D), mainly due to electron-electron interactions39. Therefore, the Co L edge of CoO3 exhibits five main features near 776.1 (A), 777.6 (B), 778.8 (C), 781.0 (D), and 793.4 (E) eV, between 772.5 and 798.5 eV, which are in agreement with earlier reports 39, 40. CoO6 and CoO8 also show similar five features. The XANE spectrum of Co3O4 shows mainly three peaks around 777.7, 779.4, and 794.0 eV, in line with an earlier report 39. The Co edge of Co3O4 has only two main features. Note that the multiplet structure of the -edge is highly sensitive to the coordination of the Co ion. The Co 2p spectrum of CoO is due to octahedral Co2+ crystal field splitting and is significantly different from Co3O4, in which both octahedral Co2+ and tetrahedral Co3+ coordinations are present.
@@
The 2p orbitals of the oxygen ligand are involved in the bonding configuration with Co metal ions. Therefore, dipole-allowed 1s to 2p transitions of O 1s can directly provide information about the oxygen charge state and reflect the partial density of O 2p unoccupied electronic states in the conduction band 41, 42. We measured the O 1s XANE spectra of CoO3, CoO6, and CoO8 along with bulk Co3O4, as shown in Fig. 3(b).The O 1s spectra of these samples exhibit mainly five features at around 529.5 (P1), 532.3 (P2), 536.5 (P3), 539.8 (P4), and 549.5 (P5) eV, ranging between 525.5 to 552.5 eV. The peaks P2 to P5 in these nanoparticles are in close agreement with earlier reports 42, 39, 40. However, an additional pre-edge feature, P1, is found compared to bulk CoO, and it becomes more pronounced with the decrease in crystallite size. This feature indicates that the electronic structure is modified in nanoparticles compared to the bulk counterpart. The features between 530 and 535 eV correspond to transitions to unoccupied electronic states of O 2p character hybridized with Co 3d states 42. Above 535 eV, features in XANE are mainly associated with transitions to O 2p character hybridized with Co 4s, 4p states 42. The O 1s character of Co3O4 is significantly different from that of CoO nanoparticles. For example, a sharp peak that appears in Co3O4 is absent in CoO nanoparticles. These results indicate that the bonding configuration of O ions with Co metal ions, and also valence states, are different in CoO nanoparticles and Co3O4. Therefore, the XANES results of CoO3, CoO6, and CoO8 clearly show that these nanoparticles exhibit a pure Co2+ valence state, which is in line with XRD data
@@ 3.3 Magnetic Properties
We have also recorded the loop at 5 K in the FC condition to probe the conventional exchange bias (Fig.6(b, c)). The sample has been cooled to 5 K from 400 K in the presence of = 7 T, and the field is cycled between 7 T, which is sufficient to obtain technical saturation and hence rules out the possibility of the minor hysteresis loop. In this case, we have recorded the loops of only CoO6 and CoO8, which show prominent hysteresis at 5 K in ZFC (Fig.5(b, c)). Under the FC condition, an enhancement in coercivity as compared to the ZFC protocol has been found in the CoO nanoparticles (Fig.6(b, c)). The hysteresis loop of CoO8 at 5 K exhibits a larger EB of around Oe in FC at 5 K. In contrast, although CoO6 does not show any exchange bias in the ZFC condition, it shows Oe in FC at 5 K. In this context, we note the reports of Lierop et al. of SEB and CEB of around 75 and 150 Oe, respectively, in NiFe/Co3O4 thin film21, and Takano et al. estimated and observed CEB of Oe on NiFe/CoO bilayers 8. The value of CEB in the CoO nanoparticles is larger compared to earlier reports on NiFe/CoO bilayers and NiFe/Co3O4 thin film 21, 8.
@@
(ii) CoO8, i.e., 10.8 nm CoO nanoparticles coated with diamagnetic silica (SiO2).
@@
When we coated CoO8 with SiO2, which is supposed to have the same crystallite size, it influences the magnetic response of nanoparticles quite interestingly. It should be mentioned here that although the thickness of the silica coating and hence the mass of the CoO/SiO2 system is not precisely known, its thickness is believed to be smaller than the diameter of the nanoparticle 34. The temperature-dependent magnetization measured in ZFC and FC at 100 Oe is shown in Fig.9(a). As temperature decreases, ZFC magnetization decreases slowly compared to CoO8 (Fig.4(c)) and shows a steep increase below 20 K (Fig. 9(a), inset). However, a clear hump associated with is seen near the same temperature of 213 K as found also in CoO8. This means SiO2 coating just affects the surface moment but not the core of the nanoparticles. The magnetization measured in FC at 7 T is approximately independent of temperature down to 20 K, and then steeply increases as temperature further decreases, similar to ZFC magnetization.
The curve at 300 K of SiO2-coated CoO nanoparticles is depicted in Fig.9(b). It shows finite coercivity and remanence, although it does not saturate till Tesla, which indicates a weak ferromagnetic response of CoO8 still present after SiO2 coating. The value of the coercive field has been found to be around Oe, which is significantly reduced compared to CoO8 (610 Oe). Note that both the and curves of SiO2-coated CoO nanoparticles show distinct behavior compared to bare CoO8, which underlines the fact that SiO2 coating affects the overall magnetic response of the CoO nanoparticles. SiO2-coating on nanoparticles is expected to reduce agglomeration of nanoparticles and likely to reduce the interparticle interaction43, 51. Thus, the SiO2-coating may affect the overall collective magnetic response. In addition, the magnetism at the interface between magnetic CoO nanoparticles and non-magnetic SiO2 can be quite interesting and is a subject of detailed investigation 52, 53.
Therefore, it is noteworthy to mention a few distinct changes in the magnetic response of Co/CoO nanocomposites and SiO2-coated CoO8 nanoparticles compared to bare CoO8 nanoparticles:
@@
(ii) The magnetization of CoO8 is significantly reduced after SiO2 coating. Since the magnetic moment in CoO8 mainly comes from surface spins, the reduction in the magnetic moment may indicate that surface spins are affected by SiO2.
(iii) The is observed near the same temperature of 213 K. This suggests that SiO2 coating mainly affects the surface moment but not the core of nanoparticles.
(iv) The bifurcation between ZFC and FC curves has reduced. This indicates that spin disorder reduces after SiO2 coating, possibly originating from the shell of nanoparticles.
(v) A steep rise in magnetization below 20 K, similar to the larger-sized sample CoO3, is observed. This is consistent with the reduction in magnetization with SiO2, similar to an increase in size as in CoO3.